Fermi National Laboratory
 

Operation Summary for March 1, 2010 - March 3, 2010

 

Monday March 1

The day shift began with Operations monitoring the Tevatron (TeV) store 7641 (185.71E30) and the Recycler stash (112.5E10), with the Antiproton Source stacking (19.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Recycler personnel made a short access into the Pelletron. Preaccelerator personnel accessed the H- Source for extractor repairs. During this time, RF technicians replaced a MI RF modulator. Water group personnel found and fixed a LCW leak on a magnet in the Meson Test beam line.

During the evening shift, the Pbar Accumulator injection kicker failed. Pbar personnel replaced a faulted high-voltage cable, which took two hours. Operations terminated store 7641 at 6:30 PM. CDF made a 1.5 hour access into their collision hall. Operations established store 7644 at 11:38 PM with an Average Initial Luminosity (AIL) of 298.6E30.

Tuesday March 2

The midnight shift began with Operations monitoring store 7644 (260.6E30) and the stash (53.2E10), with Pbar stacking (11.2E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 7:14 AM, CDF’s luminosity reading went to zero; they reported having a problem with their CLC detector and did not take data for about 22 minutes.

During the day shift, Operations terminated store 7644 at 12:03 PM. TeV personnel performed a proton only study. RF personnel repaired a MI RF modulator.

During the evening shift, TeV personnel conducted studies. Operations established store 7647 at 6:24 PM with an AIL of 303.1E30. A Booster RF station tripped off on a LCW flow fault; it reset.

Wednesday March 3

The midnight shift began with Operations monitoring store 7647 (134.75E30) and the stash (160.6E10), with Pbar stacking (30.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operators added about 1.5 inches of water to the LRF5 RF water-skid.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Meson MTest experiment T953 will start taking beam this afternoon; they will run for seven days.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/03/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory